# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v3

@@ 3.3 Complexity Analysis of Approximate Tensor Product
In this section, we analyze the computational complexity of the proposed approximate tensor product using CP decomposition. We first investigate the rank of the tensor product. Let $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ denote the CP rank of the three-way tensor ${\bm{M}}\in\mathbb{R}^{d_{3}\times d_{1}\times d_{2}}$, representing the minimal rank $R$ such that there is an equality for Eq.˜2, i.e.
@@
By using CP decomposition, the computational cost for evaluating the approximate tensor product in Eq.˜3 reduces to $\mathcal{O}\bigl(R(d_{1}+d_{2}+d_{3})\bigr)$, a significant reduction compared to the full tensor contraction cost of $\mathcal{O}(d_{1}d_{2}d_{3})$ using Eq.˜1. Similarly, the parameter budget decreases from $\mathcal{O}(d_{1}d_{2}d_{3})$ to $\mathcal{O}(R(d_{1}+d_{2}+d_{3}))$, which provides substantial savings when $R$ is small.
@@ 3.4 Error Bound and Universality Analysis
Given a rank $R\leq\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$, one seeks CP-decomposition-based approximation $\widehat{\bm{M}}$ by minimizing $\lVert{\bm{M}}-\widehat{\bm{M}}\rVert$ in a chosen norm, typically the Frobenius norm. Although the non‑convex optimization may admit spurious local minima, the optimization of CP decomposition possesses an essentially unique best rank‑$R$ approximation under Kruskal’s condition (kolda2009tensor, ), and modern alternating least‑squares or gradient methods converge to it under mild coherence assumptions (yang2023global, ). For error estimation in our setting, we specialize to the CG tensor product, where we let all irreps have the same maximum degree and all dimensions equal such that $d=d_{1}=d_{2}=d_{3}$. Given the singular values $\sigma_{k}^{(n)}$ of the mode-$n$ matricization ${\bm{M}}_{(n)}$ of $\bm{M}$} and truncating each ${\bm{M}}_{(n)}$ to rank $R_{T}=\lceil R^{1/3}\rceil$, a priori approximation error bound (de2000multilinear, ) gives
$\bigl\lVert{\bm{M}}-\widehat{{\bm{M}}}\bigr\rVert_{F}\;\leq\;\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2}.$
@@
Theorem 3.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
@@
Theorem 3.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ if $R$ is sufficiently large.
@@ 4 Experiments
The effectiveness of our method is assessed on two benchmarks: the PubChemQCR dataset, which contains millions of molecular‑dynamics snapshots, and the established Open Catalyst Project (OCP) datasets. Section˜4.1 describes both benchmarks and model configurations. Section˜4.2 compares our model with several baselines on PubChemQCR and its subset PubChemQCR-S, while Section˜4.3 and Section˜4.4 report results of our model compared to several tensor‑product baselines on OC20 and OC22 from OCP, respectively. Section˜4.5 then analyses the computational efficiency of our approach as the maximum angular degree in the tensor product is varied. Section˜4.6 further reports ablations of the proposed components across our model and a second widely used architecture.
@@ 4.1 Experimental Setup
Figure 3: Scaling behaviours of the CP-decomposition-based tensor product under different maximum angular degree scheduler: (a) approximation error, (b) $\rm{SO}(3)$-equivariance error, and (c) runtime. The error and runtime of the CP decomposition-based tensor product depend on the chosen rank, multiplicities, and the maximum angular degree, and are not tied to a specific dataset.
@@
We approximate the fully connected CG tensor product using the rank-$R$ CP decomposition across various configurations of the maximum angular degree $L$. Four rank schedules are evaluated: $R=(L+1)^{4}$, $R=16\log^{2}(L+1)$, $R=7L$, and $R=7L^{2}$. For each configuration we measure (i) Approximation Error: The relative CP tensor product error, calculated as $\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})-\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}/\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}$; (ii) Equivariance Error: The expected $\mathrm{SO}(3)$-equivariance error, defined as $\mathbb{E}_{{\bm{R}},{\bm{x}},{\bm{y}}}\bigl[\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\bigr]$, where the expectation is averaged over $1000$ random rotations $\bm{R}$} and random vectors ${\bm{x}},{\bm{y}}$; and (iii) Execution Time: The runtime for the CP-decomposition-based tensor product under each rank schedule. Our results demonstrate that the logarithmic schedule $R=7L^{2}$ consistently achieves the lowest approximation and equivariance errors across all $L$, while requiring significantly less computation time compared to the CG tensor product baseline. Thus, this schedule provides the best balance between accuracy and computational cost and is adopted for all subsequent experiments. Further rank ablations are described in Appendix˜D.
@@
Building on the capabilities of graph transformers, we develop the Tensor-Decomposition Network (TDN) by modifying the publicly available Equiformer architecture (liao2022equiformer, ). Specifically, we replace every channel-wise linear projection, normalization layer, and activation in Equiformer with batched matrix operations that act simultaneously on the multiplicity dimension of each irrep block, eliminating the need for costly slicing and reshaping between tensor and vector representations. Additionally, the depth-wise tensor product mechanism is removed and the core CG tensor products within the self-attention mechanism are substituted with our rank-$R$ CP-decomposition-based tensor product from Section˜3.1, integrated with the path-weight-sharing scheme from Section˜3.2. This design preserves the expressive power of CG tensor product while significantly improving memory and computational efficiency. We further integrate the equivariant linear layer with the path-weight-sharing scheme. For the OCP dataset, the same model architecture is applied and a subset of experiments additionally incorporate initial node embeddings following (qu2024the, ). The detailed model configurations for TDN are described in Appendix˜C.
@@ 4.2 Results on PubChemQCR
We evaluate our method on both PubChemQCR‑S and the full PubChemQCR dataset. For the small split we compare against nine state‑of‑the‑art models: SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full PubChemQCR dataset, we benchmark against SchNet and PaiNN, the only baselines that scale to its size within our hardware budget. Performance is reported as mean absolute error (MAE) for energies and root‑mean‑square error (RMSE) for forces over the validation and testing splits. The results are shown in Table˜1, the proposed TDN model achieves the lowest energy and force prediction errors across PubChemQCR‑S and the full PubChemQCR dataset, outperforming all baseline methods. In addition, the performance of TDN improves further as the size of the dataset increases.
@@ 4.3 Results on OC20
Training Setup. For the IS2RE-Direct task, we follow Equiformer’s optimization setup (liao2022equiformer, ) by an AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-3}$, a batch size of 32, and a cosine-decay learning-rate schedule. A warm-up is employed for 2 epochs on IS2RE-Direct, with a warm-up factor of 0.2; the cosine decay then runs over 30 training epochs. IS2RE-Direct experiments are run on a single NVIDIA RTX A6000-48GB GPU.
@@ 4.4 Results on OC22
Training Setup. On OC22 IS2RE, we use the same optimization setup as OC20 by an AdamW optimizer with a learning rate of $2\times 10^{-4}$, a weight decay of $10^{-3}$, a batch size of 32, and a cosine learning-rate decay with a 2-epoch warm-up for 1000 epochs on a single NVIDIA A100-80GB GPU.
@@ 4.5 Efficiency of CP-Decomposition-Based Tensor Product
Model Throughput (samples/sec) Parameter
@@
Since TDN is derived from Equiformer by replacing each CG block with CP decomposition and incorporating path-weight sharing, and removing depth-wise tensor product mechanism, we further benchmark end-to-end throughput and parameter count for both networks on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128; detailed model configurations are provided in Appendix˜D. The results in Table˜4 show that TDN processes $2.47\times$ to $8.44\times$ more structures per second and uses $63\%$$92\%$ fewer parameters than Equiformer. As the maximum degree $L$ increases, TDN achieves higher throughput and requires fewer parameters. A comprehensive time ablation of CP decomposition and path-weight sharing mechanism is provided in Table˜8.
4.6 Ablation Study of CP-Decomposition-Based Tensor Product
Table 5: Ablation study of the path-weight sharing tensor product (PS) and CP-decomposition-based tensor product (CP). Results are shown for TDN and NequIP (batzner20223, ) on PubChemQCR-S dataset.
Model Validation
Energy MAE Force RMSE Training Time
(eV) $\downarrow$ (eV/Å) $\downarrow$ (min/epoch) $\downarrow$
TDN w/o CP + PS 0.01274 0.09274 19.0
TDN 0.01295 0.09653 4.2 ($\times$ 0.22)
NequIP 0.01122 0.09009 7.5
NequIP + CP + PS 0.01115 0.09249 2.0 ($\times$ 0.26)
To further evaluate the performance and efficiency contributions of our proposed component, we perform ablations on the PubChemQCR-S dataset by removing the path-weight sharing tensor product and the CP decomposition of TDN while holding all other components fixed. As shown in Table˜5, starting from TDN, disabling both mechanisms yields slightly lower validation energy MAE and force RMSE while reducing training time roughly 78%, indicating only a minor effect on predictive accuracy. TDN is trained with a maximum degree $L=2$, four graph transformer layers, a hidden dimension of 64, and for 60 epochs. We also evaluate NequIP (batzner20223, ) by augmenting the base model with path-weight sharing tensor product and CP decomposition. This preserves downstream accuracy while cutting computational cost nearly 74%. NequIP is trained for 100 epochs with a maximum degree $L=2$, four interaction blocks, and multiplicity 64. All experiments are conducted on a single NVIDIA RTX A6000-48GB GPU.
@@ 5 Limitations
Our approach accelerates $\rm{SO}(3)$‑equivariant tensor products by replacing the full CG tensor product with a low‑rank CP decomposition. We substantiate both its theoretical speed‑up and its empirical efficacy across multiple benchmarks. One limitation of our approach is rank selection: the minimal CP rank needed to attain a given approximation error is unknown in general, and computing it exactly is NP-hard (kolda2009tensor, ). We therefore employ an empirical rank scheduler; however, this scheduler is tailored to the CG tensor and must be re-derived when extending to other equivariant tensor products. A second limitation is that path-weight sharing mechanism, while reducing parameters and memory, can introduce a small accuracy drop. In future work, we will (i) study broader families of group-equivariant tensor products, e.g., $\rm{SO}(2)$ convolution (passaro2023reducing, ) and Gaunt tensor products (luoenabling, ), to characterize their optimal rank profiles and develop general rank-selection criteria and adaptive rank selector, and (ii) design adaptive path-selection and grouping strategies that preserve the efficiency benefits of weight sharing while recovering performance.

### v4

@@ 3.3 Complexity Analysis of Approximate Tensor Product
In this section, we analyze the computational complexity of the proposed approximate tensor product using CP decomposition. We first investigate the rank of the tensor product. Let $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ denote the CP rank of the three-way tensor ${\bm{M}}\in\mathbb{R}^{d_{3}\times d_{1}\times d_{2}}$, representing the minimal rank $R$ such that there is an equality for Eq.˜3, i.e.
@@
By using CP decomposition, the computational cost for evaluating the approximate tensor product in Eq.˜4 reduces to $\mathcal{O}\bigl(R(d_{1}+d_{2}+d_{3})\bigr)$, a significant reduction compared to the full tensor contraction cost of $\mathcal{O}(d_{1}d_{2}d_{3})$ using Eq.˜2. Similarly, the space complexity decreases from $\mathcal{O}(d_{1}d_{2}d_{3})$ to $\mathcal{O}(R(d_{1}+d_{2}+d_{3}))$, which provides substantial savings when $R$ is small.
@@ 3.4 Error Bound and Universality Analysis
Given a rank $R\leq\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$, one seeks CP-decomposition-based approximation $\widehat{\bm{M}}$ by minimizing $\lVert{\bm{M}}-\widehat{\bm{M}}\rVert$ in a chosen norm, typically the Frobenius norm. Although the non‑convex optimization may admit spurious local minima, the optimization of CP decomposition possesses an essentially unique rank‑$R$ approximation under Kruskal’s condition (kolda2009tensor, ), and modern alternating least‑squares or gradient methods converge to it under mild coherence assumptions (yang2023global, ). For error estimation in our setting, we specialize to the CG tensor product, where we let all irreps have the same maximum degree and all dimensions equal such that $d=d_{1}=d_{2}=d_{3}$. Given the singular values $\sigma_{k}^{(n)}$ of the mode-$n$ matricization ${\bm{M}}_{(n)}$ of $\bm{M}$} and truncating each ${\bm{M}}_{(n)}$ to rank $R_{T}=\lfloor R^{1/3}\rfloor$, a priori approximation error bound (de2000multilinear, ) gives
$\bigl\lVert{\bm{M}}-{\bm{M}}_{\text{truncated}}\bigr\rVert_{F}\;\leq\;\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2}.$
@@
Theorem 3.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded representations ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lfloor R^{1/3}\rfloor$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
@@
Theorem 3.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-representations ${\bm{x}}\in V_{1}\cong\mathbb{R}^{d_{1}}$, ${\bm{y}}\in V_{2}\cong\mathbb{R}^{d_{2}}$ and co-domain $V_{3}\cong\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ with $R\leq d_{1}d_{2}$.
@@ 4 Experiments
The effectiveness of our method is assessed on two benchmarks: the PubChemQCR dataset, which contains millions of molecular relaxation snapshots, and the established Open Catalyst Project (OCP) datasets. Section˜4.1 describes both benchmarks and model configurations. Section˜4.2 compares our model with several baselines on PubChemQCR and its subset PubChemQCR-S, while Section˜4.3 and Section˜4.4 report results of our model compared to several tensor‑product baselines on OC20 and OC22 from OCP, respectively. Section˜4.5 then analyses the computational efficiency of our approach as the maximum angular degree in the tensor product is varied. Section˜4.6 further reports ablations of the proposed components across our model and a second widely used architecture.
@@ 4.1 Experimental Setup
Figure 3: Scaling behaviours of the CP-decomposition-based tensor product under different maximum angular degree schedules: (a) approximation error, (b) $\rm{SO}(3)$-equivariance error, and (c) runtime. The error and runtime of the CP decomposition-based tensor product depend on the chosen rank, multiplicities, and the maximum angular degree, and are not tied to a specific dataset.
@@
We approximate the fully connected CG tensor product using the rank-$R$ CP decomposition across various configurations of the maximum angular degree $L$. Four rank schedules are evaluated: $R=(L+1)^{4}$, $R=16\log^{2}(L+1)$, $R=7L$, and $R=7L^{2}$. For each configuration we measure (i) Approximation Error: The relative CP tensor product error, calculated as $\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})-\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}/\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}$; (ii) Equivariance Error: The expected $\mathrm{SO}(3)$-equivariance error, defined as $\mathbb{E}_{{\bm{R}},{\bm{x}},{\bm{y}}}\bigl[\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\bigr]$, where the expectation is averaged over $1000$ random rotations $\bm{R}$} and random vectors ${\bm{x}},{\bm{y}}$; and (iii) Execution Time: The runtime for the CP-decomposition-based tensor product under each rank schedule. Among the tested schedules, the quadratic schedule $R=7L^{2}$ consistently achieves the best accuracy-efficiency trade-off, while also requiring significantly less computation time compared to the CG tensor product baseline. Thus, this schedule provides the best balance between accuracy and computational cost and is adopted for all subsequent experiments. Further rank ablations are described in Appendix˜D.
@@
Building on the capabilities of graph transformers, we develop the Tensor-Decomposition Network (TDN) by modifying the publicly available Equiformer architecture (liao2022equiformer, ). Specifically, we replace every channel-wise linear projection, normalization layer, and activation in Equiformer with batched matrix operations that act simultaneously on the multiplicity dimension of each irrep block, eliminating the need for costly slicing and reshaping between tensor and vector representations. Additionally, the depth-wise tensor product mechanism is removed and the core CG tensor products within the self-attention mechanism are substituted with our rank-$R$ CP-decomposition-based tensor product from Section˜3.1, integrated with the path-weight-sharing scheme from Section˜3.2. This design preserves the expressive power of CG tensor product while significantly improving memory and computational efficiency. Note that for each maximum degree $L$, we precompute a rank-$R$ CP decomposition of the CG coefficient tensor once and cache the factors. These factors are treated as constants during all training runs. We further integrate the equivariant linear layer with the path-weight-sharing scheme. For the OCP dataset, the same model architecture is applied and a subset of experiments additionally incorporate initial node embeddings following (qu2024the, ). The detailed model configurations for TDN are described in Appendix˜C.
@@ 4.2 Results on PubChemQCR
We evaluate our method on both PubChemQCR‑S and the full PubChemQCR dataset. For the small split we compare against nine state‑of‑the‑art models: SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full PubChemQCR dataset, we benchmark against SchNet and PaiNN, the only baselines that scale to its size within our hardware budget. Performance is reported as mean absolute error (MAE) for energies and root‑mean‑square error (RMSE) for forces over the validation and testing splits. Note that all results are selected based on the lowest validation energy error. The results are shown in Table˜1, the proposed TDN model achieves the lowest energy and force prediction errors across PubChemQCR‑S and the full PubChemQCR dataset, outperforming all baseline methods. In addition, the performance of TDN improves further as the size of the dataset increases.
@@ 4.3 Results on OC20
Training Setup. For the IS2RE-Direct task, we follow Equiformer’s optimization setup (liao2022equiformer, ) by a maximum angular degree $L=1$, an AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-3}$, a batch size of 32, and a cosine-decay learning-rate schedule. A warm-up is employed for 2 epochs on IS2RE-Direct, with a warm-up factor of 0.2; the cosine decay then runs over 30 training epochs. IS2RE-Direct experiments are run on a single NVIDIA RTX A6000-48GB GPU.
@@ 4.4 Results on OC22
Training Setup. On OC22 IS2RE, we use the same optimization setup as OC20 by a maximum angular degree $L=1$, an AdamW optimizer with a learning rate of $2\times 10^{-4}$, a weight decay of $10^{-3}$, a batch size of 32, and a cosine learning-rate decay with a 2-epoch warm-up for 1000 epochs on a single NVIDIA A100-80GB GPU.
@@ 4.5 Efficiency of CP-Decomposition-Based Tensor Product
Model Throughput (samples/sec) Parameters
@@
Since TDN is derived from Equiformer by replacing each CG block with CP decomposition and incorporating path-weight sharing, and removing depth-wise tensor product mechanism, we further benchmark end-to-end throughput and parameter count for both networks on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128; detailed model configurations are provided in Appendix˜D. The results in Table˜4 show that TDN processes $2.47\times$ to $8.44\times$ more structures per second and uses $63\%$$92\%$ fewer parameters than Equiformer. As the maximum degree $L$ increases, TDN achieves higher throughput and requires fewer parameters. A comprehensive time ablation of CP decomposition and path-weight sharing mechanism is provided in Table˜7.
4.6 Efficiency Ablation of CP-Decomposition-Based Tensor Product
Table 5: Efficiency ablation of the path-weight sharing tensor product (PS) and CP-decomposition-based tensor product (CP). Results are shown for TDN and NequIP (batzner20223, ) on PubChemQCR-S dataset.
Model Training Time (min/epoch)
TDN w/o CP + PS 19.0
TDN 4.2 ($\times$ 0.22)
NequIP 7.5
NequIP + CP +PS 2.0 ($\times$ 0.26)
To further evaluate the efficiency contributions of our proposed components, we perform ablations on the PubChemQCR-S dataset by removing the path-weight sharing tensor product and the CP decomposition of TDN while holding all other components fixed. As shown in Table˜5, starting from TDN, enabling both mechanisms reduces training time by roughly 78%. TDN is trained with a maximum degree $L=2$, four graph transformer layers, a hidden dimension of 64. We also evaluate NequIP (batzner20223, ) by augmenting the base model with path-weight sharing tensor product and CP decomposition. This cuts computational cost nearly 74%. NequIP is trained with a maximum degree $L=2$, four interaction blocks, and multiplicity 64. All experiments are conducted on a single NVIDIA RTX A6000-48GB GPU. A full table of ablation studies including accuracy results is presented in Table˜9.
@@ 5 Limitations
Our approach accelerates $\rm{SO}(3)$‑equivariant tensor products by replacing the full CG tensor product with a low‑rank CP decomposition. We substantiate both its theoretical speed‑up and its empirical efficacy across multiple benchmarks. One limitation of our approach is rank selection: the minimal CP rank needed to attain a given approximation error is unknown in general, and computing it exactly is NP-hard (kolda2009tensor, ). We therefore employ an empirical rank scheduler; however, this scheduler is tailored to the CG tensor and must be re-derived when extending to other equivariant tensor products. A second limitation is that path-weight sharing mechanism, while reducing parameters and memory, can introduce a small accuracy drop. In future work, we will (i) study broader families of group-equivariant tensor products, e.g., $\rm{SO}(2)$ convolution (passaro2023reducing, ), Gaunt tensor products (luoenabling, ), or more complex tensor products (xie2025price, ), to characterize their optimal rank profiles and develop general rank-selection criteria and adaptive rank selector, and (ii) design adaptive path-selection and grouping strategies that preserve the efficiency benefits of weight sharing while recovering performance.